# CoverUp: Coverage-Guided LLM-Based Test Generation Juan Altmayer Pizzorno

### v1

@@ VII Conclusion
[8]
K. Sen, D. Marinov, and G. Agha, “Cute: a concolic unit testing engine for
C,” in Proceedings of the 10th European Software Engineering
Conference Held Jointly with 13th ACM SIGSOFT International Symposium on
Foundations of Software Engineering, ser. ESEC/FSE-13. New York, NY, USA: Association for Computing Machinery,
2005, p. 263–272. [Online]. Available:
@@
[10]
G. Fraser and A. Zeller, “Mutation-driven generation of unit tests and
oracles,” IEEE Transactions on Software Engineering, vol. 38, no. 2,
pp. 278–292, 2012.
[11]
G. Fraser and A. Arcuri, “Evosuite: automatic test suite generation for
object-oriented software,” in Proceedings of the 19th ACM SIGSOFT
Symposium and the 13th European Conference on Foundations of Software
Engineering, ser. ESEC/FSE ’11. New
York, NY, USA: Association for Computing Machinery, 2011, p. 416–419.
[Online]. Available: https://doi.org/10.1145/2025113.2025179
[12]
A. Panichella, F. M. Kifetew, and P. Tonella, “Reformulating branch coverage
as a many-objective optimization problem,” in 2015 IEEE 8th
(ICST), 2015, pp. 1–10.
[13]
——, “Automated test case generation as a many-objective optimisation
problem with dynamic selection of the targets,” IEEE Transactions on
Software Engineering, vol. 44, no. 2, pp. 122–158, 2018.
[14]
M. Tufano, D. Drain, A. Svyatkovskiy, S. K. Deng, and N. Sundaresan, “Unit
test case generation with transformers and focal context,” 2021.
[15]
J. Wang, Y. Huang, C. Chen, Z. Liu, S. Wang, and Q. Wang, “Software testing
with large language models: Survey, landscape, and vision,” 2024.
[16]
M. Schäfer, S. Nadi, A. Eghbali, and F. Tip, “An empirical evaluation of
using large language models for automated unit test generation,” IEEE
Transactions on Software Engineering, vol. 50, no. 1, pp. 85–105, 2024.
[17]
P. Bareiß, B. Souza, M. d’Amorim, and M. Pradel, “Code generation tools
(almost) for free? A study of few-shot, pre-trained language models on
code,” 2022.
[18]
K. Claessen and J. Hughes, “Quickcheck: a lightweight tool for random testing
of haskell programs,” in Proceedings of the Fifth ACM SIGPLAN
International Conference on Functional Programming, ser. ICFP ’00. New York, NY, USA: Association for Computing
Machinery, 2000, p. 268–279. [Online]. Available:
https://doi.org/10.1145/351240.351266
[19]
V. Vikram, C. Lemieux, and R. Padhye, “Can large language models write good
property-based tests?” 2023.
[20]
S. K. Lahiri, S. Fakhoury, A. Naik, G. Sakkas, S. Chakraborty, M. Musuvathi,
P. Choudhury, C. von Veh, J. P. Inala, C. Wang, and J. Gao, “Interactive
code generation via test-driven user-intent formalization,” 2023.
[21]
C. S. Xia, M. Paltenghi, J. Le Tian, M. Pradel, and L. Zhang, “Fuzz4all:
Universal fuzzing with large language models,” Proc. IEEE/ACM ICSE,
@@
[22]
J. A. Pizzorno and E. D. Berger, “Slipcover: Near Zero-Overhead Code
Coverage for Python,” in Proceedings of ACM SIGSOFT International
Symposium on Software Testing and Analysis (ISSTA 2023), 2023.
[23]
J. White, Q. Fu, S. Hays, M. Sandborn, C. Olea, H. Gilbert, A. Elnashar,
J. Spencer-Smith, and D. C. Schmidt, “A prompt pattern catalog to enhance
prompt engineering with ChatGPT,” 2023.
[24]
A. Gyori, A. Shi, F. Hariri, and D. Marinov, “Reliable testing: detecting
state-polluting tests to prevent test dependency,” in Proceedings of
the 2015 International Symposium on Software Testing and Analysis, ser.
ISSTA 2015. New York, NY, USA:
Association for Computing Machinery, 2015, p. 223–233. [Online]. Available:
https://doi.org/10.1145/2771783.2771793
[25]
G. Fraser and A. Arcuri, “A large-scale evaluation of automated unit test
generation using EvoSuite,” ACM Trans. Softw. Eng. Methodol.,
vol. 24, no. 2, dec 2014. [Online]. Available:
https://doi-org.silk.library.umass.edu/10.1145/2685612
[26]
A. Zeller, “Yesterday, my program worked. Today, it does not. Why?”
SIGSOFT Softw. Eng. Notes, vol. 24, no. 6, p. 253–267, oct 1999.
[Online]. Available: https://doi.org/10.1145/318774.318946
[27]
O. Parry, G. M. Kapfhammer, M. Hilton, and P. McMinn, “A survey of flaky
tests,” ACM Trans. Softw. Eng. Methodol., vol. 31, no. 1, oct 2021.
[Online]. Available: https://doi.org/10.1145/3476105
[28]
W. Lam, S. Winter, A. Wei, T. Xie, D. Marinov, and J. Bell, “A large-scale
longitudinal study of flaky tests,” Proc. ACM Program. Lang., vol. 4,
no. OOPSLA, nov 2020. [Online]. Available:
https://doi.org/10.1145/3428270
[29]
J. Turner, “New directions in communications (or which way to the information
age?),” Comm. Mag., vol. 24, no. 10, p. 8–15, oct 1986. [Online].
Available: https://doi.org/10.1109/MCOM.1986.1092946
[30]
R. Widyasari, S. Q. Sim, C. Lok, H. Qi, J. Phan, Q. Tay, C. Tan, F. Wee, J. E.
Tan, Y. Yieh, B. Goh, F. Thung, H. J. Kang, T. Hoang, D. Lo, and E. L. Ouh,
“BugsInPy: a database of existing bugs in Python programs to enable
controlled testing and debugging studies,” in Proceedings of the 28th
ACM Joint Meeting on European Software Engineering Conference and Symposium
on the Foundations of Software Engineering, ser. ESEC/FSE ’20. ACM, Nov. 2020. [Online]. Available:
http://dx.doi.org/10.1145/3368089.3417943
[31]
S. Lukasczyk, F. Kroiß, and G. Fraser, “An empirical study of automated
unit test generation for python,” CoRR, vol. abs/2111.05003, 2021.
[Online]. Available: https://arxiv.org/abs/2111.05003
[33]
N. Li, X. Meng, J. Offutt, and L. Deng, “Is bytecode instrumentation as good
as source code instrumentation: An empirical study with industrial tools
(experience report),” in IEEE 24th International Symposium on
Software Reliability Engineering, ISSRE 2013, Pasadena, CA, USA, November
4-7, 2013. IEEE Computer Society,
2013, pp. 380–389. [Online]. Available:
https://doi.org/10.1109/ISSRE.2013.6698891
[34]
R. DeMillo, R. Lipton, and F. Sayward, “Hints on test data selection: Help for
the practicing programmer,” Computer, vol. 11, no. 4, pp. 34–41,
1978.

### v2

@@ A Large-Scale Evaluation of Automated Unit Test
Gruber et al. (2024)
Martin Gruber,
Muhammad Firhard Roslan, Owain Parry,
Fabian Scharnböck, Phil McMinn, and
Gordon Fraser. 2024.
Do Automatic Test Generation Tools Generate Flaky
Tests?. In Proceedings of the IEEE/ACM 46th
International Conference on Software Engineering (Lisbon, Portugal)
(ICSE ’24). Association for
Computing Machinery, New York, NY, USA, Article
47, 12 pages.
https://doi.org/10.1145/3597503.3608138
Gyori et al. (2015)
Alex Gyori, August Shi,
Farah Hariri, and Darko Marinov.
2015.
Reliable testing: detecting state-polluting tests
to prevent test dependency. In Proceedings of the
2015 International Symposium on Software Testing and Analysis (Baltimore,
MD, USA) (ISSTA 2015).
Association for Computing Machinery,
New York, NY, USA, 223–233.
https://doi.org/10.1145/2771783.2771793
Hałas (2019)
Konrad Hałas.
2019.
MutPy, a mutation testing tool for
Python.
https://github.com/mutpy/mutpyRetrieved on 2024-09-11.
Lahiri et al. (2022)
Shuvendu K. Lahiri,
Aaditya Naik, Georgios Sakkas,
Piali Choudhury, Curtis von Veh,
Madanlal Musuvathi, Jeevana Priya Inala,
Chenglong Wang, and Jianfeng Gao.
2022.
Interactive Code Generation via Test-Driven
User-Intent Formalization.
CoRR abs/2208.05950
(2022).
https://doi.org/10.48550/ARXIV.2208.05950
arXiv:2208.05950
Lam et al. (2020)
Wing Lam, Stefan Winter,
Anjiang Wei, Tao Xie,
Darko Marinov, and Jonathan Bell.
2020.
A large-scale longitudinal study of flaky tests.
Proc. ACM Program. Lang.
4, OOPSLA, Article
202 (November 2020),
29 pages.
https://doi.org/10.1145/3428270
Lemieux et al. (2023)
Caroline Lemieux,
Jeevana Priya Inala, Shuvendu K. Lahiri,
and Siddhartha Sen. 2023.
CodaMosa: Escaping Coverage Plateaus in Test
Generation with Pre-trained Large Language Models. In
2023 IEEE/ACM 45th International Conference on
Software Engineering (ICSE). 919–931.
https://doi.org/10.1109/ICSE48619.2023.00085
Lukasczyk and Fraser (2022)
Stephan Lukasczyk and
Gordon Fraser. 2022.
Pynguin: Automated Unit Test Generation for
Python. In 44th IEEE/ACM International
Conference on Software Engineering: Companion Proceedings, ICSE Companion
2022, Pittsburgh, PA, USA, May 22-24, 2022.
ACM/IEEE, 168–172.
https://doi.org/10.1145/3510454.3516829
Lukasczyk et al. (2023)
Stephan Lukasczyk, Florian
Kroiß, and Gordon Fraser.
2023.
An empirical study of automated unit test
generation for Python.
Empir. Softw. Eng. 28,
2 (2023), 36.
https://doi.org/10.1007/S10664-022-10248-W
Miller et al. (1990)
Barton P. Miller, Lars
Fredriksen, and Bryan So.
1990.
An Empirical Study of the Reliability of UNIX
Utilities.
Commun. ACM 33,
12 (1990), 32–44.
https://doi.org/10.1145/96267.96279
Pacheco et al. (2007)
Carlos Pacheco,
Shuvendu K. Lahiri, Michael D. Ernst,
and Thomas Ball. 2007.
Feedback-Directed Random Test Generation. In
29th International Conference on Software
Engineering (ICSE 2007), Minneapolis, MN, USA, May 20-26, 2007.
IEEE Computer Society, 75–84.
https://doi.org/10.1109/ICSE.2007.37
Panichella et al. (2015)
Annibale Panichella,
Fitsum Meshesha Kifetew, and Paolo
Tonella. 2015.
Reformulating Branch Coverage as a Many-Objective
Optimization Problem. In 2015 IEEE 8th
(ICST). 1–10.
https://doi.org/10.1109/ICST.2015.7102604
Panichella et al. (2018)
Annibale Panichella,
Fitsum Meshesha Kifetew, and Paolo
Tonella. 2018.
Automated Test Case Generation as a Many-Objective
Optimisation Problem with Dynamic Selection of the Targets.
IEEE Transactions on Software Engineering
44, 2 (2018),
122–158.
https://doi.org/10.1109/TSE.2017.2663435
Parry et al. (2021)
Owain Parry, Gregory M.
Kapfhammer, Michael Hilton, and Phil
McMinn. 2021.
A Survey of Flaky Tests.
ACM Trans. Softw. Eng. Methodol.
31, 1, Article 17
(October 2021), 74 pages.
https://doi.org/10.1145/3476105
Rosas et al. (2024)
Miguel Romero Rosas,
Miguel Torres Sanchez, and Rudolf
Eigenmann. 2024.
Should AI Optimize Your Code? A Comparative Study of
Current Large Language Models Versus Classical Optimizing Compilers.
arXiv:2406.12146 [cs.AI]
https://arxiv.org/abs/2406.12146
Ryan et al. (2024)
Gabriel Ryan, Siddhartha
Jain, Mingyue Shang, Shiqi Wang,
Xiaofei Ma, Murali Krishna Ramanathan,
and Baishakhi Ray. 2024.
Code-Aware Prompting: A Study of Coverage-Guided
Test Generation in Regression Setting using LLM.
Proc. ACM Softw. Eng.
1, FSE (2024),
951–971.
https://doi.org/10.1145/3643769
Schäfer et al. (2024)
Max Schäfer, Sarah Nadi,
Aryaz Eghbali, and Frank Tip.
2024.
An Empirical Evaluation of Using Large Language
Models for Automated Unit Test Generation.
IEEE Transactions on Software Engineering
50, 1 (2024),
85–105.
https://doi.org/10.1109/TSE.2023.3334955
Sen et al. (2005)
Koushik Sen, Darko
Marinov, and Gul Agha. 2005.
CUTE: a concolic unit testing engine for C. In
Proceedings of the 10th European Software
Engineering Conference Held Jointly with 13th ACM SIGSOFT International
Symposium on Foundations of Software Engineering (Lisbon, Portugal)
(ESEC/FSE-13). Association for
Computing Machinery, New York, NY, USA,
263–272.
@@
Tillmann and de Halleux (2008)
Nikolai Tillmann and
Jonathan de Halleux. 2008.
Pex-White Box Test Generation for .NET. In
Tests and Proofs - 2nd International Conference,
TAP 2008, Prato, Italy, April 9-11, 2008. Proceedings(Lecture Notes in Computer Science,
Vol. 4966), Bernhard
Beckert and Reiner Hähnle (Eds.).
Springer, 134–153.
https://doi.org/10.1007/978-3-540-79124-9_10
Tufano et al. (2020)
Michele Tufano, Dawn
Drain, Alexey Svyatkovskiy, Shao Kun
Deng, and Neel Sundaresan.
2020.
Unit Test Case Generation with Transformers.
CoRR abs/2009.05617
(2020).
arXiv:2009.05617
https://arxiv.org/abs/2009.05617
Turner (1986)
J. Turner.
1986.
New directions in communications (or which way to
the information age?).
Comm. Mag. 24,
10 (October 1986),
8–15.
https://doi.org/10.1109/MCOM.1986.1092946
Vikram et al. (2023)
Vasudev Vikram, Caroline
Lemieux, and Rohan Padhye.
2023.
Can Large Language Models Write Good Property-Based
Tests?
CoRR abs/2307.04346
(2023).
https://doi.org/10.48550/ARXIV.2307.04346
arXiv:2307.04346
Wang et al. (2024)
Junjie Wang, Yuchao
Huang, Chunyang Chen, Zhe Liu,
Song Wang, and Qing Wang.
@@
Software Testing With Large Language Models:
Survey, Landscape, and Vision.
IEEE Trans. Software Eng.
50, 4 (2024),
911–936.
https://doi.org/10.1109/TSE.2024.3368208
White et al. (2023)
Jules White, Quchen Fu,
Sam Hays, Michael Sandborn,
Carlos Olea, Henry Gilbert,
Ashraf Elnashar, Jesse Spencer-Smith,
and Douglas C. Schmidt. 2023.
A Prompt Pattern Catalog to Enhance Prompt
Engineering with ChatGPT.
arXiv:2302.11382 [cs.SE]
Widyasari et al. (2020)
Ratnadira Widyasari,
Sheng Qin Sim, Camellia Lok,
Haodi Qi, Jack Phan,
Qijin Tay, Constance Tan,
Fiona Wee, Jodie Ethelda Tan,
Yuheng Yieh, Brian Goh,
Ferdian Thung, Hong Jin Kang,
Thong Hoang, David Lo, and
Eng Lieh Ouh. 2020.
BugsInPy: a database of existing bugs in Python
programs to enable controlled testing and debugging studies. In
ESEC/FSE ’20: 28th ACM Joint European Software
Engineering Conference and Symposium on the Foundations of Software
Engineering, Virtual Event, USA, November 8-13, 2020,
Prem Devanbu, Myra B.
Cohen, and Thomas Zimmermann (Eds.).
ACM, 1556–1560.
https://doi.org/10.1145/3368089.3417943
Xia et al. (2024)
Chunqiu Steven Xia, Matteo
Paltenghi, Jia Le Tian, Michael Pradel,
and Lingming Zhang. 2024.
Fuzz4All: Universal Fuzzing with Large Language
Models. In Proceedings of the IEEE/ACM 46th
International Conference on Software Engineering (Lisbon, Portugal)
(ICSE ’24). Association for
Computing Machinery, New York, NY, USA, Article
126, 13 pages.
https://doi.org/10.1145/3597503.3639121